OpenStudy (imnotgoodatmath):

Which literal equation is equivalent to a = 4b + 5

OpenStudy (imnotgoodatmath):

hold on, sorry

OpenStudy (mayankdevnani):

\[\huge \bf a=4b+5\] \[\huge \bf 4b=a-5\] \[\huge \bf b=\frac{a-5}{4}\] The above equation is equivalent to a=4b+5

OpenStudy (mayankdevnani):

to prove,just take value of a =1 in your equation, \[\huge \bf a=4b+5\] \[\huge \bf a=4+5=9\] Now, put b=1 in my equation, \[\huge \bf 1=\frac{a-5}{4}\] \[\huge \bf a-5=4\] \[\huge \bf a=9\]

OpenStudy (mayankdevnani):

hence,proved.

OpenStudy (mayankdevnani):

understood? @ImNotGoodAtMath

OpenStudy (imnotgoodatmath):

oh, ok hey are ou good at civics? (goverment)

OpenStudy (mayankdevnani):

i don't think so,but i will try

OpenStudy (imnotgoodatmath):

ok, ill add you hold on